Q3 Balance of Payments (I)

The citizens of Shortland, whose currency is the squat, conducted the following transactions in 2008/09:

Item Billions of squats
Exports of goods and services 40
Imports of goods and services 50
Net income and transfer x
Borrowing from the rest of the world 2
Lending to the rest of the world 12
Change in reserve assets 0

(a) What is the value of net income and transfers (x)?

(b) Set out the balance of payments accounts for Shortland.

(c) Does Shortland have a floating exchange rate?
Answer #1

Answer Balance따 fayhont val of net incone and transfos C): = .42-62 net income and tra nger (x) =-2。 -ile value alon ce oRyn

ament account Balance = yo-Jo tal accort balance =d-12 Cpheedto maintain yclource ape射 to detomire He amend Pte o aredl under ked exchange red Undr -HeechaJe tn raa ve, auels 8. and